# Offline Model-Based Optimization: Comprehensive Review

### v1

@@ 4 Surrogate Modeling
As shown in Eq. (2), we typically employ the mean squared error loss to fit the surrogate model—a loss that can also be interpreted as a maximum likelihood loss when accounting for uncertainty (Chen et al., 2024). In this work, we decompose surrogate modeling into four key components: input representation, model type, training strategy, and sampling strategy.
@@ 4.1 Input Representation
A less common discrete representation is permutation encoding, which is used when the relative order of elements is crucial. This encoding is typical in problems such as the traveling salesman problem (Lust & Teghem, 2010) and the capacitated vehicle routing problem (Zajac & Huber, 2021b). However, these problems generally do not rely on surrogate models since their final solutions can be evaluated efficiently.
@@ 4.2 Model Type
Neural networks are a class of parametric models where the architecture (e.g., the number of layers and neurons per layer) is predetermined prior to training (Salakhutdinov, 2014). These models approximate the target function by optimizing a fixed set of parameters, making them well-suited for high-dimensional and complex tasks. Owing to their generalization capabilities and scalability, NN-based surrogate models have become increasingly popular in offline MBO (Trabucco et al., 2022).
Some approaches use data-agnostic neural networks—such as multi-layer perceptrons (MLPs)—as surrogate models (Trabucco et al., 2021; Yu et al., 2021; Yuan et al., 2023; Chen et al., 2023a). While these methods demonstrate the overall effectiveness of the optimization framework, simple MLPs may struggle to accurately capture the intricacies of complex black-box functions, especially when the design itself contains rich semantic information. To better exploit domain-specific information, recent works have designed specialized architectures tailored to the data. For example, Lee et al. (2023) employs a graph neural network (GNN) designed for modeling molecular properties and guiding molecule generation, while Chen et al. (2025) uses a protein language model to extract residue embeddings that feed into a GNN for predicting antibody binding affinity, thereby steering antibody structures towards a more stable conformation.
@@
In offline MBO, the offline dataset is often limited, making non-parametric kernel-based models particularly attractive. Their effective complexity increases with the number of data points, enabling them to capture intricate function behaviors while providing principled uncertainty estimates (Rasmussen & Williams, 2006). Although offline MBO does not allow further queries of the black-box function, uncertainty quantification remains crucial: it helps identify regions in the design space where the surrogate’s predictions are less reliable. This information can be used to guide conservative optimization strategies that avoid over-optimistic predictions in poorly sampled areas.
Chen et al. (2022b) demonstrates that the neural tangent kernel—associated with infinite-width neural networks—can be more effective than standard kernels like the RBF kernel. Furthermore, Chen et al. (2023b) introduces a kernel parameterized by pre-trained biological language models for biological sequence design. In cases where pre-trained models are unavailable, deep kernel learning can be employed to learn the kernel directly from the data, with scalability achieved via inducing points (Wilson et al., 2016).
@@ 4.3 Training Strategy
Auxiliary losses are incorporated into surrogate models to refine the training process by targeting specific model behaviors, particularly regularization. For instance, Trabucco et al. (2021) trains the surrogate to systematically underestimate the true objective on out-of-distribution inputs by identifying potential adversarial examples via gradient ascent and penalizing the surrogate’s predictions at these points, thereby enforcing a conservative estimate. Similarly, Yu et al. (2021) smooths the training data with a Gaussian filter, finds the weight perturbation that maximizes the loss, and then adjusts the model parameters to maintain local smoothness with respect to both the inputs and the weights. In the same vein, Dao et al. (2024a; b) incorporate measures of model sharpness and sensitivity, respectively, as additional regularizers.
Further, Chen et al. (2022b; 2023b) propose a bidirectional learning framework that integrates knowledge from an offline dataset into a high-performing design using both forward and backward loss mappings. Here, a kernel-based model provides a closed-form solution while the backward mapping serves as a regularizer to mitigate the impact of out-of-distribution inputs. Strictly speaking, this framework does not regularize the surrogate itself but rather the behavior of the optimized designs.
In Qi et al. (2022a), offline MBO is reframed as a domain adaptation problem by treating the offline dataset as the source domain and the optimized designs as the target domain. A distance loss between these distributions is employed to train the surrogate, ensuring it produces mediocre predictions when the optimized designs deviate significantly from the training data. Furthermore, Hoang et al. (2024) shows that the quality of optimization correlates with how well the surrogate aligns with the latent gradient field underlying the offline data, and it proposes a loss function to enforce this gradient matching. Moreover, ranking losses have been shown to offer greater robustness compared to the traditional pointwise mean squared error loss: Chen et al. (2023a) employs a pairwise ranking loss, whereas Tan et al. (2025) utilizes a listwise ranking loss. Finally, auxiliary losses naturally extend to multi-objective settings. Multi-task learning techniques such as GradNorm (Chen et al., 2018) and PcGrad (Yu et al., 2020) can be used so that learning one property via a surrogate task aids in predicting another—a valuable capability in domains like biology where labeled data is often limited (Xu et al., 2022).
@@
Surrogate models often guide the sampling process of generative models, and in turn, several studies leverage insights from generative models to further enhance surrogate modeling. For instance, Fannjiang & Listgarten (2020) employs a variational autoencoder to model both the offline data distribution and the search model distribution, using the ratio of these probabilities as an importance weight to retrain the surrogate. Similarly, Qi et al. (2022a) trains a GAN discriminator to differentiate between the offline distribution and the desired design distribution, subsequently fine-tuning the surrogate to yield mediocre predictions when designs deviate significantly from the offline data. Moreover, Chen et al. (2024) derives a conditional distribution from a diffusion model, which is used to regularize the surrogate by minimizing the KL divergence between the surrogate’s output and the derived distribution.
@@ 4.4 Sampling Strategy
Selecting an optimal learning rate $\eta$ poses a significant challenge in offline MBO due to the absence of a dedicated validation set. While Beckham et al. (2024) suggests the introduction of a validation set in offline MBO, this method may require sacrificing some high-performance data. On the other hand, Trabucco et al. (2022) recommends a learning rate of $2\sqrt{d}$ for discrete tasks and $0.05\sqrt{d}$ for continuous tasks, where $d$ represents the dimension of the design space. The higher learning rate in discrete spaces accommodates their unique properties, and scaling by $\sqrt{d}$ compensates for the increase in the overall gradient norm as dimensionality increases. Additionally, Chen et al. (2023b) proposes training an auxiliary model to provide weak supervision signals for optimizing the learning rate, thus enhancing sampling effectiveness. Similarly, Chemingui et al. (2024) formulates offline MBO as an offline reinforcement learning problem, where a learned policy takes the current design as input and outputs the optimal learning rate; however, this approach may still be vulnerable to reward hacking due to out-of-distribution issues in the surrogate model.
@@
The number of optimization steps, denoted as $T$, is another vital hyperparameter. Determining the appropriate $T$ is challenging due to the absence of a ground-truth function, which raises concerns about overfitting. This hyperparameter also correlates with the learning rate: a higher learning rate might necessitate a smaller $T$ to avoid deviating from the distribution. The strategy suggested by Trabucco et al. (2021) involves using $50$ steps to generate adversarial samples and regularizing the surrogate model based on these samples. Subsequently, a similar $50$-step approach is employed during the sampling phase. Meanwhile, Yu et al. (2021); Fu & Levine (2021) report that their methods remain robust even as $T$ increases. Furthermore, Damani et al. (2023) proposes training a binary classifier to distinguish offline data from design data, observing that the degree of distribution shift depends on $T$. The classifier logits serve as a proxy for distribution shift, allowing users to constrain $T$ to regions where the surrogate predictions remain reliable.
@@
A common sampling strategy involves adapting the surrogate model at the current optimization point. While this concept overlaps with the training strategies discussed in Section 4.3, the focus here is on fine-tuning the surrogate locally. Given the impracticality of optimizing the surrogate globally, it is more feasible to refine its performance near the current point. For example, Fu & Levine (2021) estimates the conditional normalized maximum likelihood by incorporating the current point into the surrogate modeling process, and (Yu et al., 2021) adjusts the surrogate to increase local smoothness. Moreover, Chen et al. (2023a); Yuan et al. (2023) generate pseudo pairwise and pointwise labels in the neighborhood to further refine the surrogate’s local behavior.
@@ 5 Generative Modeling
In addition to surrogate modeling described in Section 4, generative modeling plays a pivotal role in offline MBO. The high-dimensionality of design spaces renders exploration challenging, and generative models offer an effective means to navigate these spaces. As detailed in Eq. (7), offline optimization can be framed as a conditional generation problem, where the objective is to model the distribution $p(\bm{x}\mid\bm{y}_{c})$ with $\bm{y}_{c}$ representing the target conditions. By Bayes’ rule, this distribution is proportional to the product of the prior $p(\bm{x})$ and the likelihood $p(\bm{y}_{c}\mid\bm{x})$. Broadly, two categories of conditional generation emerge in this context:
@@ 5.1 Variational Autoencoder (VAE)
[…]
and introduces an approximate posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$ to facilitate efficient inference.
In particular, the VAE adopts an encoder–decoder architecture: the encoder approximates the posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$, while the decoder models the likelihood $p_{\bm{\theta}}(\bm{x}|\bm{z})$. The model is trained by maximizing the evidence lower bound (ELBO):
$\mathcal{L}_{\mathrm{VAE}}(\bm{x})=\mathbb{E}_{q_{\bm{\psi}}(\bm{z}|\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}|\bm{z})\right]-\mathrm{KL}\Bigl{(}q_{\bm{\psi}}(\bm{z}|\bm{x})\,\|\,p(\bm{z})\Bigr{)}.$
Both the likelihood term $p_{\bm{\theta}}(\bm{x}|\bm{z})$ and the prior $p(\bm{z})$ (typically chosen as $\mathcal{N}(\bm{0},\bm{I})$) are usually modeled as Gaussian distributions. The KL divergence acts as a regularizer that promotes stable training, though balancing reconstruction accuracy with latent regularization—and addressing issues like posterior collapse—often requires careful tuning (Higgins et al., 2017). Sampling from a VAE is performed by drawing a latent code $\bm{z}$ from the prior $p(\bm{z})$ and then generating $\bm{x}$ using the decoder $p_{\bm{\theta}}(\bm{x}|\bm{z})$.
@@
In contrast, forward methods are more commonly employed, leveraging the continuous latent space of the VAE. Here, an unconditional VAE is first trained to embed designs into a continuous latent space, after which a surrogate model supplies gradient information to steer the latent codes toward regions corresponding to improved designs. This approach is particularly beneficial for discrete design optimization, as demonstrated by works on general discrete designs (Yu et al., 2021) and discrete molecules (Gómez-Bombarelli et al., 2018), which map these designs into a latent space amenable to gradient-based optimization. Together, these methods illustrate how VAEs can effectively navigate high-dimensional design spaces, either by directly conditioning on target attributes via the inverse method or by leveraging gradient-driven manipulations within the latent space via the forward method.
We also briefly compare VAEs and normalizing flows (Kobyzev et al., 2020), as both models map designs to latent spaces and back. While VAEs rely on a learned encoder–decoder architecture, normalizing flows use carefully designed invertible operators to establish a one-to-one correspondence between the latent and input spaces. Lee et al. (2025) observes that this invertibility effectively mitigates the reconstruction gap often seen in VAEs—which can cause property discrepancies between original and reconstructed designs. Consequently, they propose a normalizing flow model for MBO, including the SeqFlow variant for sequence designs, to address these issues directly. Although the application of normalizing flows in offline model-based optimization (MBO) remains relatively limited, they represent a promising direction for future research.
@@ 5.2 Generative Adversarial Network (GAN)
Originally proposed for image generation, GANs are renowned for their ability to produce sharper and more realistic outputs compared to VAEs, although they are also known for training instability. Various strategies, such as Wasserstein GAN (Arjovsky et al., 2017), have been introduced to improve convergence. For sampling, GANs draw noise and use the generator to synthesize realistic samples. Generally speaking, unlike VAEs, standard GANs do not provide an explicit latent code for a given sample, though extensions like BiGANs (Donahue et al., 2017) offer a means to recover latent representations.
@@
GAN-based conditional generation is typically achieved via inverse methods, owing to the absence of an inherent latent code in standard GANs. In such frameworks, both the generator and discriminator are conditioned on target labels to steer the sampling process. For instance, Kumar & Levine (2020) employs a conditional GAN where the discriminator, parameterized as $D_{\bm{\psi}}(\mathbf{x}\mid y)$, is trained to output $1$ for valid $(\mathbf{x},y)$ pairs (i.e., when $\mathbf{x}$ comes from the data and $y=f(\mathbf{x})$) and 0 for generated pairs $\bigl{(}G_{\bm{\theta}}(\mathbf{z},y),\,y\bigr{)}$. Here, the generator acts as the inverse mapping $G_{\bm{\theta}}(\mathbf{z},y)$, taking both the latent noise $\mathbf{z}$ and the condition label $y$ as inputs. This setup is optimized using the following objective:
$\min_{\bm{\theta}}\max_{\bm{\psi}}\;\mathcal{L}_{p}(\mathcal{D})=\mathbb{E}_{y\sim p(y)}\Biggl{[}\mathbb{E}_{\mathbf{x}\sim p_{\mathcal{D}}(\mathbf{x}\mid y)}\Bigl{[}\log D_{\bm{\psi}}(\mathbf{x}\mid y)\Bigr{]}+\mathbb{E}_{\mathbf{z}\sim p_{0}(\mathbf{z})}\Bigl{[}\log\Bigl{(}1-D_{\bm{\psi}}\bigl{(}G_{\bm{\theta}}(\mathbf{z},y)\mid y\bigr{)}\Bigr{)}\Bigr{]}\Biggr{]}.$ (9)
This formulation corresponds to matching the true conditional distribution $p_{\mathcal{D}}(\mathbf{x}\mid y)$ with the model distribution $p_{G_{\bm{\theta}}}(\mathbf{x}\mid y)$ (obtained by marginalizing over $\mathbf{z}$). During guided sampling, for a given target $y$, the latent variable $\mathbf{z}$ is first sampled and then optimized such that the output of the forward model, $f_{\bm{\phi}}\bigl{(}G_{\bm{\theta}}(\mathbf{z},y)\bigr{)}$, closely approximates $y$. This process quantifies the agreement between the learned inverse map and an independently trained forward model $f_{\bm{\phi}}$, ensuring that the generated sample $G_{\bm{\theta}}(\mathbf{z},y)$ not only satisfies the desired condition but also lies on the valid data manifold. Importance reweighting is also employed to construct a $p(y)$ that assigns high probability to high $y$ values.
Besides these inverse methods, the discriminator of GANs is often utilized to detect whether a design is out-of-distribution, thereby regulating the optimization process of surrogate models (Qi et al., 2022b; Yao et al., 2024). Although both surrogate models and generative models have been explored in this context, these methods are technically not GAN-based forward methods, as they do not optimize the latent code within the GAN framework but rather use the discriminator to regularize the behavior of surrogate models.

### v2

@@ 4 Surrogate Modeling
Offline MBO learns a surrogate model that approximates the black-box oracle and leverages it to guide design optimization. As shown in Eq. (2), the mean squared error loss is employed to fit the surrogate model—a loss that can also be interpreted as a maximum likelihood loss when accounting for uncertainty (Chen et al., 2024). In this work, we decompose surrogate modeling into four key components: input representation, model type, training strategy, and sampling strategy.
@@ 4.1 Input Representation
A less common discrete representation is permutation encoding, which is used when the relative order of elements is crucial. This encoding is typical in problems such as the traveling salesman problem (Lust & Teghem, 2010). However, these problems generally do not rely on surrogate models since their final solutions can be evaluated efficiently.
@@ 4.2 Model Type
Neural networks are a class of parametric models where the architecture (e.g., the number of layers and neurons per layer) is predetermined prior to training (LeCun et al., 2015). These models approximate the target function by optimizing a fixed set of parameters, making them well-suited for high-dimensional and complex tasks. Owing to their generalization capabilities and scalability, NN-based surrogate models have become increasingly popular in offline MBO (Trabucco et al., 2022).
Some approaches use data-agnostic neural networks—such as multi-layer perceptrons (MLPs)—as surrogate models (Trabucco et al., 2021; Yu et al., 2021a; Yuan et al., 2023; Chen et al., 2023a). While these methods demonstrate the overall effectiveness of the optimization framework, simple MLPs may struggle to accurately capture the intricacies of complex black-box functions, especially when the design itself contains rich semantic information. To better exploit domain-specific information, recent works have designed specialized architectures tailored to the data. For example, Lee et al. (2023) employs a graph neural network (GNN) designed for modeling molecular properties and guiding molecule generation, while Chen et al. (2025) uses a protein language model to extract residue embeddings that feed into a GNN for predicting antibody binding affinity, thereby steering antibody structures towards a more stable conformation. OmniPred (Song et al., 2024) employs a specialized neural network—specifically, a language model—trained on heterogeneous offline data comprising both textual and numerical modalities, and estimates prediction uncertainty by analyzing the concentration of sampled outputs.
@@
In offline MBO, the offline dataset is often limited, making non-parametric kernel-based models particularly attractive. Their effective complexity increases with the number of data points, enabling them to capture intricate function behaviors while providing principled uncertainty estimates (Rasmussen & Williams, 2006). Gaussian Processes are the most widely used kernel-based models, offering closed-form posterior inference and calibrated uncertainty (Frazier, 2018). Although offline MBO does not allow further queries of the black-box function, uncertainty quantification remains crucial: it helps identify regions in the design space where the surrogate’s predictions are less reliable. This information can be used to guide conservative optimization strategies that avoid over-optimistic predictions in poorly sampled areas.
Chen et al. (2022b) demonstrates that the neural tangent kernel—associated with infinite-width neural networks—can be more effective than standard kernels like the radial basis function (RBF) kernel. Furthermore, Chen et al. (2023b) introduces a kernel parameterized by pre-trained biological language models for biological sequence design. In cases where pre-trained models are unavailable, deep kernel learning can be employed to learn the kernel directly from the data, with scalability achieved via inducing points (Wilson et al., 2016).
@@ 4.3 Training Strategy
Auxiliary losses are incorporated into surrogate models to refine the training process by encouraging specific model behaviors, as detailed below.
Conservatism is a widely adopted strategy in offline MBO. Inspired by similar ideas in offline RL (Yu et al., 2021b), Trabucco et al. (2021) trains the surrogate to systematically underestimate the true objective on out-of-distribution inputs by identifying potential adversarial examples via gradient ascent and penalizing the surrogate’s predictions at these points, thereby enforcing a conservative estimate. In Qi et al. (2022), offline MBO is reframed as a domain adaptation problem by treating the offline dataset as the source domain and the optimized designs as the target domain. A distributional distance loss is used to ensure that the surrogate produces conservative (i.e., less overconfident) predictions when evaluated on samples far from the offline distribution. Further, Chen et al. (2022b; 2023b) propose a bidirectional learning framework that integrates knowledge from an offline dataset into a high-performing design using both forward and backward loss mappings. Here, a kernel-based model provides a closed-form solution while the backward mapping serves as a regularizer to mitigate the impact of out-of-distribution inputs. While this method technically regularizes the designs rather than the surrogate itself, it still promotes conservatism by encouraging generated designs to remain close to the offline dataset. This principle is closely related to distributionally robust optimization (Delage & Ye, 2010), which aims to avoid overconfident decisions under uncertain region.
Smoothness is another important strategy for improving generalization. Yu et al. (2021a) smooths the offline data with a Gaussian filter, finds the weight perturbation that maximizes the loss, and then adjusts the model parameters to maintain local smoothness with respect to both the inputs and the weights. In the same vein, Dao et al. (2024a; b) incorporate measures of model sharpness and sensitivity, respectively, to constrain the surrogate’s local behavior and improve its stability.
Ranking Consistency is another strategy used in offline MBO. In many cases, the surrogate is tasked with selecting promising designs rather than predicting precise objective values. To support this, Tan et al. (2025) employs a listwise ranking loss that encourages the surrogate to preserve the relative ordering of candidate designs. In scenarios where pointwise labels are noisy or unreliable—such as with augmented data—Chen et al. (2023a) adopts a pairwise ranking loss, demonstrating improved robustness over traditional mean squared error losses. In essence, ranking-based losses provide discrete, relative supervision that enhances the surrogate’s robustness and decision-making reliability.
Gradient Alignment is exemplified by Hoang et al. (2024), which demonstrates that the effectiveness of optimization is strongly correlated with how well the surrogate’s gradient field aligns with the underlying gradients of the offline data. To this end, the authors propose a dedicated loss function to enforce such alignment and improve optimization quality.
Finally, auxiliary losses naturally extend to multi-objective settings. Multi-task learning techniques such as GradNorm (Chen et al., 2018) and PcGrad (Yu et al., 2020) can be used so that learning one property via a surrogate task aids in predicting another—a valuable capability in domains like biology where labeled data is often limited (Xu et al., 2022).
@@
Surrogate models often guide the sampling process of generative models, and in turn, several studies leverage insights from generative models to further enhance surrogate modeling. For instance, Fannjiang & Listgarten (2020) employs a variational autoencoder to model both the offline data distribution and the search model distribution, using the ratio of these probabilities as an importance weight to retrain the surrogate. Similarly, Qi et al. (2022) trains a GAN discriminator to differentiate between the offline distribution and the desired design distribution, subsequently fine-tuning the surrogate to yield mediocre predictions when designs deviate significantly from the offline data. Moreover, Chen et al. (2024) derives a conditional distribution from a diffusion model, which is used to regularize the surrogate by minimizing the KL divergence between the surrogate’s output and the derived distribution.
Practical Guide on Selecting Training Strategies
We have discussed various training strategies, which typically complement rather than conflict with each other. Choosing the appropriate strategy depends primarily on the intended use-case and the operational context of the surrogate model. Below, we provide practical guidance for selecting among these strategies:
  • When using the surrogate model for gradient-based optimization (e.g. gradient ascent), conservatism-based approaches such as conservative objective models (COMs) (Trabucco et al., 2021) are particularly effective. For scenarios where the surrogate’s primary role is ranking and selecting promising designs, strategies emphasizing ranking consistency are crucial (Tan et al., 2025; Chen et al., 2023a). Additionally, if the surrogate exhibits sensitivity to perturbations in inputs or weights, smoothness strategies can be prioritized (Yu et al., 2021a).
  • In situations where a robust generative model is available, one may (1) employ sample reweighting, adjusting the sample weights based on the probability ratio between the offline and search distributions, thus refining surrogate training; (2) identify out-of-distribution samples by leveraging a GAN discriminator (Qi et al., 2022) or by directly computing likelihood scores (Chen et al., 2024); (3) generate additional synthetic samples with pseudo-labeling to improve the surrogate.
  • Domain knowledge injection is highly effective when pre-trained, domain-specific models are accessible. These can serve as feature extractors to improve extrapolation performance (Chen et al., 2023b).
  • Ensemble-based predictions generally outperform individual surrogate predictions, making collaborative ensembling effective when computational resources allow (Chen et al., 2023a; Yuan et al., 2023).
@@ 4.4 Sampling Strategy
Selecting an optimal learning rate $\eta$ poses a significant challenge in offline MBO due to the absence of a dedicated validation set. While Beckham et al. (2024) suggests the introduction of a validation set in offline MBO, this method may require sacrificing some high-performing data. Additionally, Chen et al. (2023b) proposes training an auxiliary model to provide weak supervision signals for optimizing the learning rate, thus enhancing sampling effectiveness. Similarly, Chemingui et al. (2024) formulates offline MBO as an offline reinforcement learning problem, where a learned policy takes the current design as input and outputs the optimal learning rate; however, this approach may still be vulnerable to reward hacking due to out-of-distribution issues in the surrogate model.
@@
The number of optimization steps, denoted as $T$, is another vital hyperparameter. Determining the appropriate $T$ is challenging due to the absence of a ground-truth function, which raises concerns about overfitting. This hyperparameter also correlates with the learning rate: a higher learning rate might necessitate a smaller $T$ to avoid deviating from the distribution. The strategy suggested by Trabucco et al. (2021) involves using $50$ steps to generate adversarial samples and regularizing the surrogate model based on these samples. Subsequently, a similar $50$-step approach is employed during the sampling phase. Meanwhile, Yu et al. (2021a); Fu & Levine (2021) report that their methods remain robust even as $T$ increases. Furthermore, Damani et al. (2023) proposes training a binary classifier to distinguish offline data from design data, observing that the degree of distribution shift depends on $T$. The classifier logits serve as a proxy for distribution shift, allowing users to constrain $T$ to regions where the surrogate predictions remain reliable.
@@
A common sampling strategy involves adapting the surrogate model at the current optimization point. While this concept overlaps with the training strategies discussed in Section 4.3, the focus here is on fine-tuning the surrogate locally. Given the impracticality of optimizing the surrogate globally, it is more feasible to refine its performance near the current point. For example, Fu & Levine (2021) estimates the conditional normalized maximum likelihood by incorporating the current point into the surrogate modeling process, and (Yu et al., 2021a) adjusts the surrogate to increase local smoothness. Moreover, Chen et al. (2023a); Yuan et al. (2023) generate pseudo pairwise and pointwise labels in the neighborhood to further refine the surrogate’s local behavior.
@@ 5 Generative Modeling
In addition to surrogate modeling described in Section 4, generative modeling plays a pivotal role in offline MBO. The high-dimensionality of design spaces renders exploration challenging, and generative models offer an effective means to navigate these spaces. As detailed in Eq. (7), offline black-box optimization can be framed as a conditional generation problem, where the objective is to model the distribution $p(\bm{x}\mid\bm{y}_{c})$ with $\bm{y}_{c}$ representing the target conditions. By Bayes’ rule, this distribution is proportional to the product of the prior $p(\bm{x})$ and the likelihood $p(\bm{y}_{c}\mid\bm{x})$. Broadly, two categories of conditional generation emerge in this context:
@@ 5.1 Variational Autoencoder (VAE)
$p(\bm{x})=\int p(\bm{x}|\bm{z})\,p(\bm{z})\,d\bm{z},$
and introduces an approximate posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$ to facilitate efficient inference. In particular, the VAE adopts an encoder–decoder architecture: the encoder approximates the posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$, while the decoder models the likelihood $p_{\bm{\theta}}(\bm{x}|\bm{z})$.
@@
In contrast, forward methods are more commonly employed, leveraging the continuous latent space of the VAE. Here, an unconditional VAE is first trained to embed designs into a continuous latent space, after which a surrogate model supplies gradient information to steer the latent codes toward regions corresponding to improved designs. This approach is particularly beneficial for discrete design optimization, as demonstrated by works on general discrete designs (Yu et al., 2021a) and discrete molecules (Gómez-Bombarelli et al., 2018), which map these designs into a latent space amenable to gradient-based optimization. Together, these methods illustrate how VAEs can effectively navigate high-dimensional design spaces, either by directly conditioning on target attributes via the inverse method or by leveraging gradient-driven manipulations within the latent space via the forward method.
We also briefly compare VAEs and normalizing flows (Kobyzev et al., 2020), as both models map designs to latent spaces and back. While VAEs rely on a learned encoder–decoder architecture, normalizing flows use carefully designed invertible operators to establish a one-to-one correspondence between the latent and input spaces. Lee et al. (2025) observes that this invertibility effectively mitigates the reconstruction gap often seen in VAEs—which can cause property discrepancies between original and reconstructed designs. Consequently, they propose a normalizing flow model for MBO, including the SeqFlow variant for sequence designs, to address these issues directly. Although the application of normalizing flows in offline MBO remains relatively limited, they represent a promising direction for future research.
@@ 5.2 Generative Adversarial Network (GAN)
GAN-based conditional generation is typically achieved via inverse methods, owing to the absence of an inherent latent code in standard GANs. In such frameworks, both the generator and discriminator are conditioned on target labels to steer the sampling process. For instance, Kumar & Levine (2020) employs a conditional GAN where the discriminator, parameterized as $D_{\bm{\psi}}(\bm{x}\mid y)$, is trained to output $1$ for valid $(\bm{x},y)$ pairs (i.e., when $\bm{x}$ comes from the data and $y=f(\bm{x})$) and 0 for generated pairs $\bigl(G_{\bm{\theta}}(\bm{z},y),\,y\bigr)$. Here, the generator acts as the inverse mapping $G_{\bm{\theta}}(\bm{z},y)$, taking both the latent noise $\bm{z}$ and the condition label $y$ as inputs. This setup is optimized using the following objective:
$\min_{\bm{\theta}}\max_{\bm{\psi}}\;\mathcal{L}_{p}(\mathcal{D})=\mathbb{E}_{y\sim p(y)}\Biggl[\mathbb{E}_{\bm{x}\sim p_{\mathcal{D}}(\bm{x}\mid y)}\Bigl[\log D_{\bm{\psi}}(\bm{x}\mid y)\Bigr]+\mathbb{E}_{\bm{z}\sim p_{0}(\bm{z})}\Bigl[\log\Bigl(1-D_{\bm{\psi}}\bigl(G_{\bm{\theta}}(\bm{z},y)\mid y\bigr)\Bigr)\Bigr]\Biggr].$ (9)
This formulation corresponds to matching the true conditional distribution $p_{\mathcal{D}}(\bm{x}\mid y)$ with the model distribution $p_{G_{\bm{\theta}}}(\bm{x}\mid y)$ (obtained by marginalizing over $\bm{z}$). During guided sampling, for a given target $y$, the latent variable $\bm{z}$ is first sampled and then optimized such that the output of the forward model, $f_{\bm{\phi}}\bigl(G_{\bm{\theta}}(\bm{z},y)\bigr)$, closely approximates $y$. This process quantifies the agreement between the learned inverse map and an independently trained forward model $f_{\bm{\phi}}$, ensuring that the generated sample $G_{\bm{\theta}}(\bm{z},y)$ not only satisfies the desired condition but also lies on the valid data manifold. Importance reweighting is also employed to construct a $p(y)$ that assigns high probability to high $y$ values.
Besides these inverse methods, the discriminator of GANs is often utilized to detect whether a design is out-of-distribution (OOD), thereby regulating the optimization process of surrogate models (Qi et al., 2022; Yao et al., 2024a). Although both surrogate models and generative models have been explored in this context, these methods are technically not GAN-based forward methods, as they do not optimize the latent code within the GAN framework but rather use the discriminator to regularize the behavior of surrogate models.